# Generation of large-scale magnetic fields upstream of gamma-ray burst afterglow shocks

### v1

@@ 4.1 Parameter Scan
To draw general conclusions regarding the impact of continuous electron-positron pair injection on magnetic field evolution, we performed a parameter scan over $\alpha$ – the final beam-to-background density ratio – and $t_{f}$ – the pair injection duration – with $\alpha$ varying between 1, 2, 4, 8, and 16 and $t_{f}\omega_{pb}$ varying between $5\times 10^{3}$, $10^{4}$, $2\times 10^{4}$, and $4\times 10^{4}$. Here, we discuss our main findings from the parameter scan.
@@
Figure 4 compares the time evolution of the magnetic energy density (left vertical axis) and the average transverse wavenumber of the magnetic field (right vertical axis) for four simulations with $\alpha=2$: $t_{f}\omega_{pb}=5\times 10^{3}$, $10^{4}$, $2\times 10^{4}$, and $4\times 10^{4}$. Across each of these cases, the trajectories of $\varepsilon_{B}$ and of $k_{y}\right>c/\omega_{pb}$ over time are qualitatively the same: early growth of primarily electrostatic modes yields a weak magnetic field, which is then amplified in energy and scale by a filamentation instability enhanced by continuous pair injection, and is finally brought to even higher energies and larger scales by a relatively slow-growing secular phase. As $t_{f}\omega_{pb}$ increases, the timescales of each of these phases also increases: the electrostatic phase lasts longer, the filamentation phase starts later and develops more slowly, and the secular phase growth is more shallow. While the decoupling of the background electrons from the background protons and of the beam electrons from the beam positrons occurs at later times for higher $t_{f}\omega_{pb}$, the final velocities of the particles do not vary significantly between simulations at fixed $\alpha$.
@@
Figure 4: Time evolution of $\varepsilon_{B}$ (solid curves, measured on the left vertical axis) and $k_{y}\right>c/\omega_{pb}$ (dashed curves, measured on the right vertical axis) for four simulations with $\alpha=2$: $t_{f}\omega_{pb}=5\times 10^{3}$ (blue), $t_{f}\omega_{pb}=10^{4}$ (orange), $t_{f}\omega_{pb}=2\times 10^{4}$ (green), and $t_{f}\omega_{pb}=4\times 10^{4}$ (pink). The green and pink curves extend well beyond the plotted time range.
@@
In agreement with Garasev & Derishev (2016), as we hold $\alpha$ fixed and increase $t_{f}\omega_{pb}$, both the final $\varepsilon_{B}$ and the final $k_{y}\right>c/\omega_{pb}$ decrease. Since smaller-scale magnetic fluctuations dissipate more quickly, these two trends are linked: a longer $t_{f}$ allows more time for smaller-scale modes to die out – thus reducing the final magnetic energy – yet also allows more time for larger-scale modes to grow. This trend towards larger scales at higher $t_{f}\omega_{pb}$ is illustrated quantitatively in Figure 5, which shows 1D transverse power spectra of $B_{z}$ at the end of pair injection for our four $\alpha=2$ cases; as $t_{f}\omega_{pb}$ increases, the peak of the spectrum shifts towards smaller $k_{y}c/\omega_{pb}$, or towards larger scales. This same trend is illustrated in Figure 6, which shows spatial profiles of $B_{z}$ at the end of injection; as $t_{f}\omega_{pb}$ increases, so does the width of the filamentary magnetic structures.
@@
Figure 6: Spatial profiles of $B_{z}/\sqrt{8\pi\left(\gamma_{b}-1\right)m_{e}n_{b}c^{2}}$ at the end of pair injection for four simulations with $\alpha=2$: (a) $t_{f}\omega_{pb}=5\times 10^{3}$, (b) $t_{f}\omega_{pb}=10^{4}$, (c) $t_{f}\omega_{pb}=2\times 10^{4}$, and (d) $t_{f}\omega_{pb}=4\times 10^{4}$. The color bar is rescaled to $\left|B_{z}\right|^{0.2}\times{\rm sign}(B_{z})$ to display a wider dynamic range.
@@
We now hold $t_{f}\omega_{pb}$ fixed and vary $\alpha$. Figure 7 shows the evolution of $\varepsilon_{B}$ and $k_{y}\right>c/\omega_{pb}$ for five simulations with $t_{f}\omega_{pb}=10^{4}$ and with $\alpha$ varying between 1, 2, 4, 8, and 16. As $\alpha$ increases, the timescales of each of the phases decrease; for the $\alpha=8$ and $\alpha=16$ cases, the secular phase saturates early enough (at $t\omega_{pb}\sim 4600$ for $\alpha=8$ and at $t\omega_{pb}\sim 3200$ for $\alpha=16$) to allow for an extended period of magnetic field decay before the end of pair injection. As $\alpha$ increases, the decoupling of the background electrons from the protons becomes more dramatic; by the end of our $\alpha=16$ simulation, the background electrons have nearly formed a single population with the beam pairs in terms of velocity and temperature. The relative drift between this lepton beam and the protons triggers a slowly-growing filamentation instability that drives the secular phase; Figure 8 shows the resulting proton filaments, which become more pronounced with higher $\alpha$.
@@
Figure 7: Time evolution of $\varepsilon_{B}$ (solid curves, measured on the left vertical axis) and $k_{y}\right>c/\omega_{pb}$ (dashed curves, measured on the right vertical axis) for five simulations with $t_{f}\omega_{pb}=10^{4}$: $\alpha=1$ (blue), $\alpha=2$ (orange), $\alpha=4$ (green), $\alpha=8$ (pink), and $\alpha=16$ (brown).
@@
Figure 8: Comparison of the ion density contrast (i.e., $n_{i}/\left<n_{i}\right>-1$, normalized to the maximum value in each panel) at saturation of the first filamentation phase and at saturation of the secular phase for the $t_{f}\omega_{pb}=10^{4}$ simulations with $\alpha=8$ (panels (a) and (b)) and $\alpha=16$ (panels (c) and (d)). The $\alpha=8$ profiles are taken at $t\omega_{pb}=950$ (panel (a)) and $t\omega_{pb}=4100$ (panel (b)); the $\alpha=16$ profiles are taken at $t\omega_{pb}=750$ (panel (c)) and $t\omega_{pb}=3000$ (panel (d)).
@@
Figure 9: Time evolution of $\sigma$ (panels (a) and (b)) and $k_{y}\right>c/\omega_{pi}$ (panels (c) and (d)) for three simulations with $\alpha/(t_{f}\omega_{pi})=2/29$ (solid curves), three with $\alpha/(t_{f}\omega_{pi})=4/29$ (dashed curves), and three with $\alpha/(t_{f}\omega_{pi})=8/29$ (dotted curves); within each group of three, the simulation with $\alpha=4$ is shown in orange, $\alpha=8$ in green, and $\alpha=16$ in pink. The horizontal axis is rescaled to $\left[\alpha/(t_{f}\omega_{pi})\right]^{1/2}t\omega_{pi}$ to capture the self-similar time evolution of $\sigma$ and $k_{y}\right>c/\omega_{pi}$. In panels (a) and (c), we show our simulation data as-is, with no vertical rescaling of $\sigma$ or $k_{y}\right>c/\omega_{pi}$; in panels (b) and (d), we rescale the normalization of $\sigma$ and $k_{y}\right>c/\omega_{pi}$ for each curve based on the pre-factors in Equations (7) and (8). This shows good agreement between our data and these analytic expressions.
@@
Considering both $\alpha$ and $t_{f}$ together, the quantity $\left[\alpha/(t_{f}\omega_{pi})\right]^{1/2}$ (for $\omega_{pi}$ the plasma frequency of background protons) seems to play a key role in determining the evolution of the magnetic field. [^2] In the following, we present scalings of various quantities with time measured as either $t\omega_{pb}$ or as $t\omega_{pi}$; the latter scales as $\sqrt{m_{e}/m_{i}}\,t\omega_{pe}$. In Figure 9, we plot the time evolution of the upstream magnetization,
@@
$\sigma\equiv\frac{\left<B_{z}^{2}\right>}{4\pi n_{i}m_{i}c^{2}}=4(\gamma_{b}-1)\alpha\frac{m_{e}}{m_{i}}\varepsilon_{B},$ (6)
@@
as well as the time evolution of $k_{y}\right>c/\omega_{pi}$ for a total of nine simulations with varying $\alpha$ and $t_{f}\omega_{pi}$: three simulations have $\alpha/(t_{f}\omega_{pi})=2/29$, three have $\alpha/(t_{f}\omega_{pi})=4/29$, and three have $\alpha/(t_{f}\omega_{pi})=8/29$. We rescale the units of time to $\left[\alpha/(t_{f}\omega_{pi})\right]^{1/2}\,t\omega_{pi}$, resulting in an identical time evolution for all the curves: the onset time of the filamentation phase, the growth rate of the filamentation phase, and the growth rate of the secular phase are all constant across the nine cases plotted, suggesting that the evolution of the magnetic field under the action of continuous pair injection exhibits remarkable self-similarity. Cases with different $\alpha/(t_{f}\omega_{pi})$ differ by an overall normalization factor; our simulation data strongly imply that
$\sigma=\left(\frac{\alpha}{t_{f}\omega_{pi}}\right)^{1/2}\,\mathcal{F}\left[t\omega_{pi}\left(\frac{\alpha}{t_{f}\omega_{pi}}\right)^{1/2}\right]$ (7)
@@
$k_{y}\right>c/\omega_{pi}=\left(\frac{\alpha}{t_{f}\omega_{pi}}\right)^{1/4}\,\mathcal{G}\left[t\omega_{pi}\left(\frac{\alpha}{t_{f}\omega_{pi}}\right)^{1/2}\right],$ (8)
@@
$\varepsilon_{B}=\frac{1}{2}\left(\frac{m_{i}}{m_{e}}\right)^{5/4}\frac{\alpha^{-1/4}}{\left(t_{f}\omega_{pb}\right)^{1/2}}\,\mathcal{F}\left[t\omega_{pb}\frac{\alpha^{1/4}}{\left(t_{f}\omega_{pb}\right)^{1/2}}\left(\frac{m_{i}}{m_{e}}\right)^{-1/4}\right]$ (9)
@@
$k_{y}\right>c/\omega_{pb}$ $=\left(\frac{m_{i}}{m_{e}}\right)^{-3/8}\frac{\alpha^{-1/8}}{\left(t_{f}\omega_{pb}\right)^{1/4}}$
×$\times\mathcal{G}\left[t\omega_{pb}\frac{\alpha^{1/4}}{\left(t_{f}\omega_{pb}\right)^{1/2}}\left(\frac{m_{i}}{m_{e}}\right)^{-1/4}\right].$ (10)
@@
$\frac{\Gamma}{\omega_{pi}}\propto\left(\frac{\alpha}{t_{f}\omega_{pi}}\right)^{1/2}$ (11)
or, in units of $\omega_{pb}$, as
$\frac{\Gamma}{\omega_{pb}}\propto\frac{\alpha^{1/4}}{\left(t_{f}\omega_{pb}\right)^{1/2}}\left(\frac{m_{i}}{m_{e}}\right)^{-1/4}.$ (12)
@@
$\frac{\Gamma}{\omega_{pi}}\propto\frac{\alpha t_{\rm on}}{t_{f}};$ (14)
@@
this implies that the instantaneous beam-to-background density ratio at the onset time of the instability sets the growth rate. The linear dependence of the growth rate on the instantaneous density contrast, as well as the fact that the dominant wavelength grows exponentially at the same rate as the magnetic field, are evocative of the nonlinear streaming instability described by Peterson et al. (2021, 2022). This linear dependence is faster than the $\alpha^{1/2}$ scaling expected in the cold-beam limit of the filamentation instability, yet slower than the $\alpha^{3/2}$ scaling expected in the ultra-relativistic, hot-beam limit (Bret et al., 2010a).
@@
While the growth rate of the filamentation phase seems to depend only on the instantaneous beam-to-background density ratio at the onset time, continuous pair injection is still crucial for sustaining this growth over an extended period of time. In Figure 10, we illustrate the importance of continuous pair injection by comparing multiple simulations (with $\alpha=2$ and $t_{f}\omega_{pb}=10^{4}$) with pair injection cut off at different times before $t_{f}$; we cut off injection at $t_{\rm cut}\omega_{pb}=1000$ (at the beginning of the filamentation phase), $t_{\rm cut}\omega_{pb}=1450$ (in the middle of the filamentation phase), $t_{\rm cut}\omega_{pb}=1750$ (immediately before the saturation of the filamentation phase), and $t_{\rm cut}\omega_{pb}=4000$ (near the beginning of the secular phase). In all cases, cutting off pair injection quenches subsequent growth of the magnetic and electric energies. Once injection is stopped, the energy briefly continues to increase (at a notably reduced growth rate) before saturating and decaying; evidently, any theoretical model of these systems must account for a growth rate with dynamical loading. This dependence on continuous injection applies to both the filamentation phase and the secular phase: the $t_{\rm cut}\omega_{pb}=1000$ and $t_{\rm cut}\omega_{pb}=1450$ cases prematurely halt the filamentation growth, while all the cut-off cases eliminate the secular growth, indicating that the secular phase is a unique element of our continuous-injection simulations.
@@
Figure 10: Comparison of simulations with $\alpha=2$ and $t_{f}\omega_{pb}=10^{4}$, but with continuous pair injection prematurely cut off at different times. The blue curves represent a simulation with fully continuous injection (i.e., an injection cutoff at $t_{\rm cut}\omega_{pb}=10^{4}$), the orange curves represent a simulation with $t_{\rm cut}\omega_{pb}=1000$, the green curves a simulation with $t_{\rm cut}\omega_{pb}=1450$, the pink curves a simulation with $t_{\rm cut}\omega_{pb}=1750$, and the brown curves a simulation with $t_{\rm cut}\omega_{pb}=4000$. The solid curves show the evolution of $\varepsilon_{B}$, the dotted curves show the evolution of $\varepsilon_{E}$, and the dashed vertical lines show the times at which injection was cut off.
@@
Continuous injection is also vital for developing and maintaining the large-scale magnetic structures discussed above. Figure 11 shows spatial profiles of the magnetic field at $t\omega_{pb}=10^{4}$ for the simulations with $t_{\rm cut}\omega_{pb}=1750$, with $t_{\rm cut}\omega_{pb}=4000$, and with no premature injection cutoff. These plots demonstrate that, in the absence of continuous injection, any large-scale magnetic structure fades away.
@@
Figure 11: Spatial profiles of $B_{z}/\sqrt{8\pi\left(\gamma_{b}-1\right)m_{e}n_{b}c^{2}}$ at $t\omega_{pb}=10^{4}$ for three simulations with $\alpha=2$ and $t_{f}\omega_{pb}=10^{4}$ but with different pair injection cutoff times: (a) $t_{\rm cut}\omega_{pb}=1750$, (b) $t_{\rm cut}\omega_{pb}=4000$, and (c) $t_{\rm cut}\omega_{pb}=10^{4}$ (i.e., no cutoff). The color bar is rescaled to $\left|B_{z}\right|^{0.2}\times{\rm sign}(B_{z})$ to display a wider dynamic range.

### v2

@@ 4.1 Parameter Scan
To draw general conclusions regarding the impact of continuous electron-positron pair injection on magnetic field evolution, we performed a parameter scan over $\alpha$ – the final beam-to-background density ratio – and $t_{\rm f}$ – the pair injection duration – with $\alpha$ varying between 1, 2, 4, 8, and 16 and $t_{\rm f}\omega_{\rm pb}$ varying between $5\times 10^{3}$, $10^{4}$, $2\times 10^{4}$, and $4\times 10^{4}$. Here, we discuss our main findings from the parameter scan.
@@
Figure 4 compares the time evolution of the magnetic energy density (left vertical axis) and the average transverse wavenumber of the magnetic field (right vertical axis) for four simulations with $\alpha=2$: $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$, $10^{4}$, $2\times 10^{4}$, and $4\times 10^{4}$. Across each of these cases, the trajectories of $\varepsilon_{B}$ and of $k_{y}\right>c/\omega_{\rm pb}$ over time are qualitatively the same: early growth of primarily electrostatic modes yields a weak magnetic field, which is then amplified in energy and scale by a filamentation instability enhanced by continuous pair injection, and is finally brought to even higher energies and larger scales by a relatively slow-growing secular phase. As $t_{\rm f}\omega_{\rm pb}$ increases, the timescales of each of these phases also increases: the electrostatic phase lasts longer, the filamentation phase starts later and develops more slowly, and the secular phase growth is more shallow. While the decoupling of the background electrons from the background protons and of the beam electrons from the beam positrons occurs at later times for higher $t_{\rm f}\omega_{\rm pb}$, the final velocities of the particles do not vary significantly between simulations at fixed $\alpha$.
@@
Figure 4: Time evolution of $\varepsilon_{B}$ (solid curves, measured on the left vertical axis) and $k_{y}\right>c/\omega_{\rm pb}$ (dashed curves, measured on the right vertical axis) for four simulations with $\alpha=2$: $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$ (blue), $t_{\rm f}\omega_{\rm pb}=10^{4}$ (orange), $t_{\rm f}\omega_{\rm pb}=2\times 10^{4}$ (green), and $t_{\rm f}\omega_{\rm pb}=4\times 10^{4}$ (pink). The green and pink curves extend well beyond the plotted time range.
@@
In agreement with Garasev & Derishev (2016), as we hold $\alpha$ fixed and increase $t_{\rm f}\omega_{\rm pb}$, both the final $\varepsilon_{B}$ and the final $k_{y}\right>c/\omega_{\rm pb}$ decrease. Since smaller-scale magnetic fluctuations dissipate more quickly, these two trends are linked: a longer $t_{\rm f}$ allows more time for smaller-scale modes to die out – thus reducing the final magnetic energy – yet also allows more time for larger-scale modes to grow. This trend towards larger scales at higher $t_{\rm f}\omega_{\rm pb}$ is illustrated quantitatively in Figure 5, which shows 1D transverse power spectra of $B_{z}$ at the end of pair injection for our four $\alpha=2$ cases; as $t_{\rm f}\omega_{\rm pb}$ increases, the peak of the spectrum shifts towards smaller $k_{y}c/\omega_{\rm pb}$, or towards larger scales. This same trend is illustrated in Figure 6, which shows spatial profiles of $B_{z}$ at the end of injection; as $t_{\rm f}\omega_{\rm pb}$ increases, so does the width of the filamentary magnetic structures.
@@
Figure 6: Spatial profiles of $B_{z}/\sqrt{8\pi\left(\gamma_{\rm b}-1\right)m_{\rm e}n_{\rm b}c^{2}}$ at the end of pair injection for four simulations with $\alpha=2$: (a) $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$, (b) $t_{\rm f}\omega_{\rm pb}=10^{4}$, (c) $t_{\rm f}\omega_{\rm pb}=2\times 10^{4}$, and (d) $t_{\rm f}\omega_{\rm pb}=4\times 10^{4}$. The color bar is rescaled to $\left|B_{z}\right|^{0.2}\times{\rm sign}(B_{z})$ to display a wider dynamic range.
@@
We now hold $t_{\rm f}\omega_{\rm pb}$ fixed and vary $\alpha$. Figure 7 shows the evolution of $\varepsilon_{B}$ and $k_{y}\right>c/\omega_{\rm pb}$ for five simulations with $t_{\rm f}\omega_{\rm pb}=10^{4}$ and with $\alpha$ varying between 1, 2, 4, 8, and 16. As $\alpha$ increases, the timescales of each of the phases decrease; for the $\alpha=8$ and $\alpha=16$ cases, the secular phase saturates early enough (at $t\omega_{\rm pb}\sim 4600$ for $\alpha=8$ and at $t\omega_{\rm pb}\sim 3200$ for $\alpha=16$) to allow for an extended period of magnetic field decay before the end of pair injection. As $\alpha$ increases, the decoupling of the background electrons from the protons becomes more dramatic; by the end of our $\alpha=16$ simulation, the background electrons have nearly formed a single population with the beam pairs in terms of velocity and temperature. The relative drift between this lepton beam and the protons triggers a slowly-growing filamentation instability that drives the secular phase; Figure 8 shows the resulting proton filaments, which become more pronounced with higher $\alpha$.
@@
Figure 7: Time evolution of $\varepsilon_{B}$ (solid curves, measured on the left vertical axis) and $k_{y}\right>c/\omega_{\rm pb}$ (dashed curves, measured on the right vertical axis) for five simulations with $t_{\rm f}\omega_{\rm pb}=10^{4}$: $\alpha=1$ (blue), $\alpha=2$ (orange), $\alpha=4$ (green), $\alpha=8$ (pink), and $\alpha=16$ (brown).
@@
Figure 8: Comparison of the ion density contrast (i.e., $n_{\rm i}/\left<n_{\rm i}\right>-1$, normalized to the maximum value in each panel) at saturation of the first filamentation phase and at saturation of the secular phase for the $t_{\rm f}\omega_{\rm pb}=10^{4}$ simulations with $\alpha=8$ (panels (a) and (b)) and $\alpha=16$ (panels (c) and (d)). The $\alpha=8$ profiles are taken at $t\omega_{\rm pb}=950$ (panel (a)) and $t\omega_{\rm pb}=4100$ (panel (b)); the $\alpha=16$ profiles are taken at $t\omega_{\rm pb}=750$ (panel (c)) and $t\omega_{\rm pb}=3000$ (panel (d)).
@@
Figure 9: Time evolution of $\sigma$ (panels (a) and (b)) and $k_{y}\right>c/\omega_{\rm pi}$ (panels (c) and (d)) for three simulations with $\alpha/(t_{\rm f}\omega_{\rm pi})=2/29$ (solid curves), three with $\alpha/(t_{\rm f}\omega_{\rm pi})=4/29$ (dashed curves), and three with $\alpha/(t_{\rm f}\omega_{\rm pi})=8/29$ (dotted curves); within each group of three, the simulation with $\alpha=4$ is shown in orange, $\alpha=8$ in green, and $\alpha=16$ in pink. The horizontal axis is rescaled to $\left[\alpha/(t_{\rm f}\omega_{\rm pi})\right]^{1/2}t\omega_{\rm pi}$ to capture the self-similar time evolution of $\sigma$ and $k_{y}\right>c/\omega_{\rm pi}$. In panels (a) and (c), we show our simulation data as-is, with no vertical rescaling of $\sigma$ or $k_{y}\right>c/\omega_{\rm pi}$; in panels (b) and (d), we rescale the normalization of $\sigma$ and $k_{y}\right>c/\omega_{\rm pi}$ for each curve based on the pre-factors in Equations (7) and (8). This shows good agreement between our data and these analytic expressions.
@@
Considering both $\alpha$ and $t_{\rm f}$ together, the quantity $\left[\alpha/(t_{\rm f}\omega_{\rm pi})\right]^{1/2}$ (for $\omega_{\rm pi}$ the plasma frequency of background protons) seems to play a key role in determining the evolution of the magnetic field. [^3] In the following, we present scalings of various quantities with time measured as either $t\omega_{\rm pb}$ or as $t\omega_{\rm pi}$; the latter scales as $\sqrt{m_{\rm e}/m_{\rm i}}\,t\omega_{\rm pe}$. In Figure 9, we plot the time evolution of the upstream magnetization,
@@
$\sigma\equiv\frac{\left<B_{z}^{2}\right>}{4\pi n_{\rm i}m_{\rm i}c^{2}}=4(\gamma_{\rm b}-1)\alpha\frac{m_{\rm e}}{m_{\rm i}}\varepsilon_{B},$ (6)
@@
as well as the time evolution of $k_{y}\right>c/\omega_{\rm pi}$ for a total of nine simulations with varying $\alpha$ and $t_{\rm f}\omega_{\rm pi}$: three simulations have $\alpha/(t_{\rm f}\omega_{\rm pi})=2/29$, three have $\alpha/(t_{\rm f}\omega_{\rm pi})=4/29$, and three have $\alpha/(t_{\rm f}\omega_{\rm pi})=8/29$. We rescale the units of time to $\left[\alpha/(t_{\rm f}\omega_{\rm pi})\right]^{1/2}\,t\omega_{\rm pi}$, resulting in an identical time evolution for all the curves: the onset time of the filamentation phase, the growth rate of the filamentation phase, and the growth rate of the secular phase are all constant across the nine cases plotted, suggesting that the evolution of the magnetic field under the action of continuous pair injection exhibits remarkable self-similarity. Cases with different $\alpha/(t_{\rm f}\omega_{\rm pi})$ differ by an overall normalization factor; our simulation data strongly imply that
$\sigma=\left(\frac{\alpha}{t_{\rm f}\omega_{\rm pi}}\right)^{1/2}\,\mathcal{F}\left[t\omega_{\rm pi}\left(\frac{\alpha}{t_{\rm f}\omega_{\rm pi}}\right)^{1/2}\right]$ (7)
@@
$k_{y}\right>c/\omega_{\rm pi}=\left(\frac{\alpha}{t_{\rm f}\omega_{\rm pi}}\right)^{1/4}\,\mathcal{G}\left[t\omega_{\rm pi}\left(\frac{\alpha}{t_{\rm f}\omega_{\rm pi}}\right)^{1/2}\right],$ (8)
@@
$\varepsilon_{B}=\frac{1}{2}\left(\frac{m_{\rm i}}{m_{\rm e}}\right)^{5/4}\frac{\alpha^{-1/4}}{\left(t_{\rm f}\omega_{\rm pb}\right)^{1/2}}\,\mathcal{F}\left[t\omega_{\rm pb}\frac{\alpha^{1/4}}{\left(t_{\rm f}\omega_{\rm pb}\right)^{1/2}}\left(\frac{m_{\rm i}}{m_{\rm e}}\right)^{-1/4}\right]$ (9)
@@
$k_{y}\right>c/\omega_{\rm pb}$ $=\left(\frac{m_{\rm i}}{m_{\rm e}}\right)^{-3/8}\frac{\alpha^{-1/8}}{\left(t_{\rm f}\omega_{\rm pb}\right)^{1/4}}$
×$\times\mathcal{G}\left[t\omega_{\rm pb}\frac{\alpha^{1/4}}{\left(t_{\rm f}\omega_{\rm pb}\right)^{1/2}}\left(\frac{m_{\rm i}}{m_{\rm e}}\right)^{-1/4}\right].$ (10)
@@
$\frac{\Gamma}{\omega_{\rm pi}}\propto\left(\frac{\alpha}{t_{\rm f}\omega_{\rm pi}}\right)^{1/2}$ (11)
or, in units of $\omega_{\rm pb}$, as
$\frac{\Gamma}{\omega_{\rm pb}}\propto\frac{\alpha^{1/4}}{\left(t_{\rm f}\omega_{\rm pb}\right)^{1/2}}\left(\frac{m_{\rm i}}{m_{\rm e}}\right)^{-1/4}.$ (12)
@@
$\frac{\Gamma}{\omega_{\rm pi}}\propto\frac{\alpha t_{\rm on}}{t_{\rm f}};$ (14)
@@
this implies that the instantaneous beam-to-background density ratio at the onset time of the instability sets the growth rate. The linear dependence of the growth rate on the instantaneous density contrast, as well as the fact that the dominant wavelength grows exponentially at the same rate as the magnetic field, are evocative of the nonlinear streaming instability described by Peterson et al. (2021, 2022). This linear dependence is faster than the $\alpha^{1/2}$ scaling expected in the cold-beam limit of the filamentation instability, yet slower than the $\alpha^{3/2}$ scaling expected in the ultra-relativistic, hot-beam limit (Bret et al., 2010).
@@
While the growth rate of the filamentation phase seems to depend only on the instantaneous beam-to-background density ratio at the onset time, continuous pair injection is still crucial for sustaining this growth over an extended period of time. In Figure 10, we illustrate the importance of continuous pair injection by comparing multiple simulations (with $\alpha=2$ and $t_{\rm f}\omega_{\rm pb}=10^{4}$) with pair injection cut off at different times before $t_{\rm f}$; we cut off injection at $t_{\rm cut}\omega_{\rm pb}=1000$ (at the beginning of the filamentation phase), $t_{\rm cut}\omega_{\rm pb}=1450$ (in the middle of the filamentation phase), $t_{\rm cut}\omega_{\rm pb}=1750$ (immediately before the saturation of the filamentation phase), and $t_{\rm cut}\omega_{\rm pb}=4000$ (near the beginning of the secular phase). In all cases, cutting off pair injection quenches subsequent growth of the magnetic and electric energies. Once injection is stopped, the energy briefly continues to increase (at a notably reduced growth rate) before saturating and decaying; evidently, any theoretical model of these systems must account for a growth rate with dynamical loading. This dependence on continuous injection applies to both the filamentation phase and the secular phase: the $t_{\rm cut}\omega_{\rm pb}=1000$ and $t_{\rm cut}\omega_{\rm pb}=1450$ cases prematurely halt the filamentation growth, while all the cut-off cases eliminate the secular growth, indicating that the secular phase is a unique element of our continuous-injection simulations.
@@
Figure 10: Comparison of simulations with $\alpha=2$ and $t_{\rm f}\omega_{\rm pb}=10^{4}$, but with continuous pair injection prematurely cut off at different times. The blue curves represent a simulation with fully continuous injection (i.e., an injection cutoff at $t_{\rm cut}\omega_{\rm pb}=10^{4}$), the orange curves represent a simulation with $t_{\rm cut}\omega_{\rm pb}=1000$, the green curves a simulation with $t_{\rm cut}\omega_{\rm pb}=1450$, the pink curves a simulation with $t_{\rm cut}\omega_{\rm pb}=1750$, and the brown curves a simulation with $t_{\rm cut}\omega_{\rm pb}=4000$. The solid curves show the evolution of $\varepsilon_{B}$, the dotted curves show the evolution of $\varepsilon_{E}$, and the dashed vertical lines show the times at which injection was cut off.
@@
Continuous injection is also vital for developing and maintaining the large-scale magnetic structures discussed above. Figure 11 shows spatial profiles of the magnetic field at $t\omega_{\rm pb}=10^{4}$ for the simulations with $t_{\rm cut}\omega_{\rm pb}=1750$, with $t_{\rm cut}\omega_{\rm pb}=4000$, and with no premature injection cutoff. These plots demonstrate that, in the absence of continuous injection, any large-scale magnetic structure fades away.
@@
Figure 11: Spatial profiles of $B_{z}/\sqrt{8\pi\left(\gamma_{\rm b}-1\right)m_{\rm e}n_{\rm b}c^{2}}$ at $t\omega_{\rm pb}=10^{4}$ for three simulations with $\alpha=2$ and $t_{\rm f}\omega_{\rm pb}=10^{4}$ but with different pair injection cutoff times: (a) $t_{\rm cut}\omega_{\rm pb}=1750$, (b) $t_{\rm cut}\omega_{\rm pb}=4000$, and (c) $t_{\rm cut}\omega_{\rm pb}=10^{4}$ (i.e., no cutoff). The color bar is rescaled to $\left|B_{z}\right|^{0.2}\times{\rm sign}(B_{z})$ to display a wider dynamic range.